CONTINUITY PRINCIPLE AND EXTENSION PROPERTIES OF MEROMORPHIC MAPPINGS WITH VALUES IN NON KÄHLER

MANIFOLDS

SERGEI IVASHKOVICH

Abstract. In this paper we are proving an analogue of E. Levi Continuity Principle for meromorphic mappings with values in general complex spaces. We also describe the singularities of meromorphic mappings into complex spaces carrying pluriclosed Hermitian metric forms and geometrical obstructions to removing them.